OpenStudy (anonymous):

WILL MEDAL & FAN !!! PLEASE HELP ASAP !!! Calculate the area of a trapezoid. The measure of the first base is 12 feet, the measure of the second is 5 feet, while the height is 2 feet. a. 14 square feet b. 15 square feet c. 16 square feet d. 17 square feet

TheSmartOne (thesmartone):

The area of a trapezoid is A=1/2 x (Base1 +base2) x height

TheSmartOne (thesmartone):

So A=.5 x (12+5) x 2

TheSmartOne (thesmartone):

A=.5 x 17 x 2

TheSmartOne (thesmartone):

A=17 feet^2

TheSmartOne (thesmartone):

Because .5 x 2 = 1

OpenStudy (anonymous):

A=a+b 2h=2+12 2·5=35ft²

OpenStudy (anonymous):

so its D ?

TheSmartOne (thesmartone):

Yup. @ninaesb

OpenStudy (anonymous):

yha its D

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

welcome